A Level Further Mathematics 9231 Paper 42, Oct/Nov 2025

Question paper and mark scheme50 marks6 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9231/42 · Oct/Nov 2025

Download PDF
1

A large company claims that the median salary of its employees is 32 500.Thesalaries() of 15 randomly selected employees are listed below. 18 750 30 500 125 000 42 500 25 000 26 000 52 500 23 000 27 500 19 500 25 500 33 000 30 000 21 500 29 000

a

Explain why a Wilcoxon signed-rank test may not be appropriate to test the company’s claim in this case.

[1]
b

Carry out a sign test at the 10% significance level to investigate the company’s claim.

[5]
2

A factory produces packets of biscuits. The total mass of biscuits in a packet has a normal distribution with mean μ. A random sample of 12 packets is taken and the mass of the contents of each packet, x g, is recorded. The results are summarised as follows. Σx=2390Σx2=476 117

a

Find a 99% confidence interval for μ.

[4]
b

A test of the null hypothesis μ=k is carried out on this sample using a 5% significance level. The test does not support the alternative hypothesis μ<k. Find the greatest possible value of k.

[3]
3

A traffic expert claims that the number of breakdowns occurring each day on a busy section of a motorway follows a Poisson distribution with mean 0.7. The number of breakdowns each day over a 200-day period was recorded. The following table contains the observed frequencies together with some of the expected frequencies using the expert’s distribution.

Number of breakdowns per day01234⩾5
Observed frequency887326733
Expected frequency99.317m24.3335.6780.994n
a

Find the value of m and the value of n, correct to 3 decimal places.

[2]
b

Carry out a goodness of fit test at the 5% significance level to investigate the expert’s claim.

[6]
4

A continuous random variable X has cumulative distribution function F given by F(x)={0x<1,15x+a1⩽x<4,150x2+b4⩽x⩽6,1x>6, where a and b are constants.

a

Find the value of a and the value of b.

[2]
b

Find the probability density function of X.

[2]
c

Given that E(X)=529150, find Var(X).

[3]
d

Find the 10th and 90th percentiles of X.

[3]
5

An engineer is comparing the tensile strengths of steel rods made from two machines, A and B. The engineer randomly selects 8 rods from machine A and 6 rods from machine B. The tensile strengths, in appropriate units, are given in the following table.

Machine A402403415412409407406410
Machine B401398395397410405

You should assume that the two distributions are normal and have the same population variance. Use a t-test at the 5% significance level to test whether there is any difference in the mean tensile strengths of steel rods from the two machines.

[9]
6

The discrete random variable X has probability generating function GX(t) given by GX(t)=t(3−2t)2.

a

Find E(X) and Var(X).

[5]
b

The discrete random variable Y has probability generating function GY(t) given by GY(t)=t2(3−2t)2. The random variable Z is the sum of the random variables X and Y. Assuming X and Y are independent, find P(Z>4).

[5]

Mark scheme

9231/42 · Oct/Nov 2025

Download PDF

1a · 1 mark

  • B1

    Data has outliers / is skewed / lacks symmetry, suggesting population/distribution may not be symmetrical.

1b · 5 marks

  • B1

    H0: population median is [$]32,500. H1: population median is not [$]32,500.

  • B1

    Test statistic = 4 (or 11)

  • B1

    [X∼B(15,0.5)] P(X≤4)=0.0592 OR P(X≥11)=0.0592

  • M1

    '0.0592' > 0.05, accept H0 / do not reject H0 / not significant.

  • A1

    Insufficient evidence to reject [company’s] claim. OR Insufficient evidence to suggest that median [salary] is not [$]32,500.

2a · 4 marks

  • M1

    s2=111(′476117′−′2390′212)[=32633=9.8788]

  • B1

    3.106

  • M1

    ′2390′12±′3.106′′9.8788′12

  • A1

    [196,202]

2b · 3 marks

  • B1

    −1.796

  • M1

    ′199.17′−k′9.8788′12>′−1.796′ or ∣′199.17′−k∣′9.8788′12<′1.796′

  • A1

    200.8

3a · 2 marks

  • B1

    m=69.522

  • B1

    n=0.156 or 0.157

3b · 6 marks

  • B1

    H0:Po(0.7) fits the data. H1:Po(0.7) does not fit the data.

  • M1

    Combining last 3 columns give 13 and 6.828 or 6.829.

  • M1

    (88−99.317)299.317+(73−69.552)269.552+⋯+(13−6.829)26.829 [1.2896+0.1740+0.1142+5.5764] OR 88299.317+73269.552+26224.333+1326.829−200

  • A1

    7.15 or 7.16

  • M1

    '7.15' < 7.815, accept H0 / do not reject H0 / not significant.

  • A1

    Insufficient evidence to reject expert’s claim. OR Insufficient evidence to suggest that Po(0.7) is not a good fit/model [for the data].

4a · 2 marks

  • B1

    a=−15

  • B1

    b=725

4b · 2 marks

  • M1

    Attempt at differentiating both parts.

  • A1

    All correct, fully defined including domain. Domain must cover all reals, must not overlap at boundaries.

4c · 3 marks

  • M1

    E(X2)=15∫14x2 dx+125∫46x3 dx=[115x3]14+[1100x4]46

  • M1

    Var(X)=735−(529150)2

  • A1

    2.16

4d · 3 marks

  • M1

    10th percentile: 15x−15=110 90th percentile: 150x2+725=910

  • A1

    10th percentile = 1.5

  • A1

    90th percentile = 31

5 · 9 marks

  • B1

    H0:μA=μBH1:μA≠μB

  • M1

    ′1331848′−′3264′28[=136]

  • M1

    ′964964′−′2406′26[=158]

  • M1

    sp2=′136′+′158′8+6−2=24.5

  • A1

    24.5

  • M1

    t = \frac{\frac{'3264'}{8} - \frac{'2406'}{6}}{\sqrt{'s_p^2' (\frac{1}{8} + \frac{1}{6})}} = 2.62

  • A1

    2.62

  • M1

    '2.62' > 2.179, reject H0 / significant.

  • A1

    Sufficient evidence to suggest that there is a difference in the mean tensile strengths of steel rods produced by the two machines.

6a · 5 marks

  • M1

    GX′(t)=(3−2t)2+4(3−2t)t(3−2t)4[=3+2t(3−2t)3]

  • A1

    E(X)=GX′(1)=5

  • M1

    GX′′(t)=2(3−2t)3+6(3+2t)(3−2t)2(3−2t)6[=24+8t(3−2t)4]

  • M1

    Var(X)=GX′′(1)+GX′(1)−(GX′(1))2=32+5−25

  • A1

    [Var(X)=]12

6b · 5 marks

  • M1

    GZ(t)=t3(3−2t)4=181t3(1+(−4)(−23t)+… )

  • M1

    P(Z=4)=their coefficient of t4. OR P(Z=3)=their coefficient of t3.

  • A1

    [P(Z=4)=]8243 and [P(Z=3)=]181

  • M1

    P(Z>4)=1−′8243′−′181′

  • A1

    P(Z>4)=232243

Questions

  1. 1
    Question 16 marks
    Mark scheme ›
  2. 2
    Question 27 marks
    Mark scheme ›
  3. 3
    Question 38 marks
    Mark scheme ›
  4. 4
    Question 410 marks
    Mark scheme ›
  5. 5
    Question 59 marks
    Mark scheme ›
  6. 6
    Question 610 marks
    Mark scheme ›

More 9231 papers

Other papers from Oct/Nov 2025

Solve it on ExamPalStart